1. The following data represent the travel time (in minutes) to school for seven students enrolled in Chaffey College. Treat the seven students as a sample.  

Student

Travel time

John

32

Jeff

24

Kevin

15

Ron

40

Noel

32

Mary

23

Christy

27

(a) Find standard deviation of the sample. (b) Find variance of the sample.                                                                                                                
Answer #1

Solution :

We are given a data of sample size n = 7

32,24,15,40,32,23,27

Using this, first we find sample mean(T) and sample standard deviation(s).

T =   sumof observations n

= (32 + 24.......+ 27)/7

T = 27.571

Now ,

s =   ΣΕ (Χ, - X)2 η – 1

s = \sqrt{}(32 - 27.571)2 + ... + (27 - 27.571)2/ 7 - 1

s = \sqrt{}385.714/6

s = 8.018

Sample standard deviation s = 8.018

The sample variance s2 ,

s2 = 8.0182 = 64.286

The sample variance s2 =  64.286